# On the handling of method failure in comparison studies

### v1

@@ 3.3.3 Check if data matches scope of study
If method failure remains after solving or ruling out method misuse, authors should assess if real-world users are likely to encounter such data (and therefore, this method failure) in practice. These considerations must include the scope of the study and are of course not purely objective. Importantly, they should be reported in the study, since they carry valuable information and affect the interpretation of the study’s results.
Consider the case of method failure caused by the method’s inability to make required calculations due to an “extreme” structure of the given simulated data set, for example highly imbalanced classes in the outcome or predictor variable(s), as in White et al., (2024). Depending on the scope of the study, these extreme data sets may be of interest or not. If they are deemed unrealistic for the considered scope and thus not relevant to the readers of the study, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic given the aim of their study, in line with White et al., (2024). Similarly, “too extreme” data may also affect real data based studies. If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely. See Section 4.2 for an illustration.
However, if the data sets causing failure are deemed relevant to the study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation. In particular, they may modify certain parameters of the concerned method(s) (see Section 3.3.4) or resort to a fallback (see Section 3.3.5).
@@ 3.3.4 Be realistic when modifying method implementations
As already touched upon, in some instances of method failure, a way of mimicking the behavior of real-world users may be to resort to an alternative implementation of the affected method (see Masaoud and Stryhn, (2010) for an example). This way, method failure may be resolved completely, allowing for the assessment of unconditional (absolute and relative) performance of the methods under investigation.
For instance, one could modify a parameter of the method, such as increasing the maximum number of iterations in the iterative maximum likelihood process or changing the internally employed optimization algorithm. However, authors of comparison studies should not tweak the implementation of the method thoughtlessly. Instead, they should always consider if the applied change in the parameter setting fits the given research stage and corresponding scope of the study. For instance, if the study goal is to understand the methods’ behaviors quite generally (and in earlier phases of research according to the terminology of Heinze et al., (2024)), authors may modify a greater variety of parameters to see when the method is successful and when it fails. However, if the study takes place in later phases and aims to guide inexperienced users, the applied parameter changes should be more restricted to what a naive user of the method would be able and likely to do. In any case, the performance assessment of the given method is only meaningful if the authors stress and justify their parameter choices when reporting the study.
Furthermore, it is important to make sure that the method comparison is still fair, meaning that the effort spent on each method under consideration should not differ significantly between the methods, especially in late phase studies, where neutrality is of crucial importance. Finally, one should be careful about increasing the runtimes of the affected method(s) when running a comparison study. Especially when the number of iterations is raised, runtime or memory consumption can increase notably and might result in runtime or memory issues so that method failure is, after all, not solved.
If, after these considerations, a change in the implementation of the given method is deemed reasonable, any deviation from the default parameter setting of a given method should be reported and justified transparently. Note that, as will be discussed in Section 3.3.5, changes in a method’s implementation could also be applied as a fallback strategy.
@@ 3.3.5 Consider a fallback strategy
When contemplating ways to mimic real-world users when encountering method failure in later research stages, Morris et al., (2019) argue that they would usually not give up but naturally resort to an alternative method, a fallback, that successfully provides an output. Therefore, it can make sense to transfer this fallback strategy into the context of methodological comparison studies and evaluate the performance of method pipelines of the type
“use method A if it successfully produces an output, use method B otherwise”.
@@
On the one hand, this approach incorporates what data analysts really do in practice, and on the other, it indirectly yields automatic handling of undefined performance values in case they affect method A but not method B. If chosen carefully and in a way that there is a suitable fallback for each instance of method failure, fallback strategies allow for the assessment of (absolute and relative) unconditional performance of the methods under investigation. Besides realistically reflecting the behavior of real-world users, this approach results in a performance assessment that depends only on the definition of the pipeline. Note that when using fallback strategies, we no longer evaluate the “pure” performance of method A. However, it is clear from Section 3.1 that this in an unachievable goal in the first place.
The concept of fallback strategies is already intuitively considered in some existing comparison studies. Voeltz, (2024), for instance, conclude from their study that “the approach by Frömke et al., (2022) could be used as a fallback strategy in case of non-convergence [of the overall best-performing method] because it always yields results”, where the approach by Frömke et al., (2022) is a non-iterative procedure, therefore never affected by non-convergence by design. Going one step further than Voeltz, (2024), one could provide a performance assessment for the resulting pipeline.
A special case of the fallback strategy has indirectly been addressed in Section 3.3.4: it may sometimes make sense to define an alternative implementation of the given method (e.g, using a different optimizer or an increased number of iterations) as a fallback instead of a completely different method (as in Crowther et al., (2014) who changed the number of quadrature points). The corresponding pipeline would then consist of the default implementation of that method for those data sets where it successfully produces an output and the alternative implementation for the remaining data sets.
In practice, the choice of a suitable fallback depends on the research phase. For instance, in early phases, the fallback may be constituted of an alternative implementation (e.g., alternative parameter choice) of the original method. Also, authors should strongly base the selection of fallbacks on what can be considered realistic given the expertise in statistics and programming of the targeted users. For instance, in a study targeting data analysts with limited statistical background, the fallback strategy may be an alternative method with default parameters rather than the same method with complex parameter settings. Since the choice of suitable fallbacks involves some subjectivity, authors should justify their choice(s) transparently.
If failures are expected at the planning stage, i.e., before conducting the study, one may consider the whole pipeline as a competing method to be assessed in the comparison study. Another approach would be to design a pipeline after seeing the results of the comparison study, as did by Voeltz, (2024). In this case, the definition of the pipeline can be viewed as a result of the comparison study rather than a method that has to be evaluated in the first place. The definition of the fallback strategy, however, may not always be as straightforward as in Voeltz, (2024). It may for example make sense to consider complex rules involving a second fallback in case the first one fails.
@@
3.3.6 Transparently report failure proportions, underlying interplays, and applied approaches
Following all recommendations above, authors should transparently report (i) any occurrences and proportions of method failure, (ii) the underlying interplay(s) causing failure (in different instances), and (iii) the corresponding approaches following the in-depth investigations of method failure.
@@ 3.3.6 Transparently report failure proportions, underlying interplays, and applied approaches
Report any occurrences and proportions of method failure  This can serve as an important criterion for real-world users selecting a suitable method. Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited. Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Voeltz, (2024) for an example. If method failure is predictable for certain settings and all affected methods fail for the same data sets (see, e.g., the example from Section 4.1), it may make sense to include the binary performance measure of the type “can handle setting: yes. vs. no” and report it alongside the corresponding number of repetitions/benchmark data sets associated with the failure.
@@
Report interplay of method and data causing failure  Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”. This enables a more comprehensive overview of the applicability of the methods compared to reporting the corresponding failure proportions only. Readers can learn in which settings a method is applicable, and no less importantly, in which it is not. Additionally, these insights serve as justification of the chosen handlings to method failure. An example of a corresponding reporting can be found in Hornung et al., (2024) who provide detailed descriptions in the supplement of their paper.
Report applied handlings  Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail. This includes the removal of any misused methods, the removal of data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and the incorporations of fallback strategies. Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
3.3.7 Finally: think ahead, pre-register but remain flexible
The previous examples and those that will be outlined in Section 4 make clear that method failure and its cause(s), related to complex dynamics between factors such as method and data set, are often difficult to anticipate. While pre-specifying the handling of undefined values has the advantage of reducing analytical uncertainty and its dangers in terms of selective reporting (Siepe et al., , 2024), pre-specifying a standard procedure prior to conducting the study may sometimes be delicate.
@@ 3.3.7 Finally: think ahead, pre-register but remain flexible
Our recommendation is, as far as possible, to consider in advance which combinations of data and methods could lead to failure. For instance, carefully reading user manuals and help pages of all methods before conducting the study might reveal specific requirements of a method on the data. This may make some instances of method failure less unexpected. More generally, making method failures less unexpected can be achieved by conducting a pilot experiment. Pilot experiments test whether an experimental design is appropriate for the specific context in which it is applied, and are recommended in several textbooks on Design and Analysis of Experiments (Cohen, , 1995; Dean et al., , 2017; Montgomery, , 2020). For example, in the context of real data comparison studies, a pilot experiment might use a fraction of the available data sets, say $20\%$, to test whether the specified method implementations produce reasonable results. The remaining $80\%$ of the data sets are then used for the main experiment, which may be refined based on what is learned in the pilot. A pilot study can also help authors anticipate possible runtime issues for which they can then pre-specify the random subset of data sets on which the experiment for the affected method will be restricted. Of course, it is important to clearly report the results of the pilot study too. After implementing these strategies, pre-specifying a procedure for handling method failure may be less hazardous. However, even despite pre-registration, researchers should be aware that unforeseen issues may appear and require adaptations of the pre-specified strategy. See Lakens, (2024) for a discussion of the consequences of deviations from pre-registration on the severity and validity of inferences.
3.4 Slimmed-down recommendations when time is short
@@
For some authors, there may be time constraints that do not allow for any (in-depth) implementations of our recommendations from the previous Section 3.3, however, authors nevertheless wish to aggregate results over data sets despite the existence of undefined values caused by method failure. In these cases, where it is difficult to follow the entirety of our recommendations, we recommend the following slimmed-down analysis and reporting strategy. Authors may present performance results after (i) discarding the corresponding data sets for the failing methods, (ii) discarding these data sets for all methods, and (iii) report the failure proportions for all methods, ideally as an additional performance measure. This three-fold analysis and reporting enables more differentiated assessments of the methods’ performances and avoids the preference of methods based solely on how data sets associated with method failure are discarded. In particular, it provides insights into how the methods successfully producing an output perform on the data sets for which others fail (and how this can influence the method comparison).
@@
However, following Section 3.1, authors should be aware and clearly state in their manuscript that any form of aggregating performance results despite existing undefined values does not allow an assessment of unconditional performance of all methods under investigation.
@@ 4 Illustrations
In the following, we illustrate the guidelines presented in Section 3 through two illustrative comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling. For both examples, we first demonstrate how the ad hoc approaches discussed in Section 3.2, such as discarding data sets associated with failure, can distort the results and even lead to misleading method recommendations. Then, we illustrate the implementation of the recommendations from Section 3.3. Thereby, our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison, rather than the method comparison itself. This has three consequences. First, both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study. Second, we do not give a comprehensive overview of all aspects of the studies but focus on those that are relevant to the issue of undefined performance values. Third, our goal is not to provide insights into the performance of the compared methods (larger dedicated studies would be necessary for this purpose).
@@ 4.1.1 General
Our first example, a statistical simulation study, deals with the comparison of odds ratio estimation methods. Odds ratios (“ORs”) quantify the association between an exposure and an outcome and are typically used in (retrospective) case-control studies. Here, we focus on the case of a binary exposure $X\in\left\{0,1\right\}$ and binary outcome $Y\in\left\{0,1\right\}$. While a straightforward way of estimating the OR is to calculate $\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$ (where $n_{ij}$ is the observed number of subjects with $X=i$ and $Y=j$), other methods employ more complex modeling strategies. Note that ORs are bounded in $(0,\infty)$, where values $>1$ imply a positive association between exposure and outcome.
With the true OR value being generally unknown in real data sets, OR estimation methods are typically compared using simulation studies under different scenarios, e.g., different sample sizes, underlying true OR values, and probabilities of exposure $p_{x}$. Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations. This results in $n_{ij}=0$ for the corresponding combination $i,j\in\{0,1\}$ for this data set. These zeros are called “sampling zeros” and cause computational issues (e.g., division by 00) for many of the available OR estimation methods. They lead to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$”, resulting in undefined performance values in the corresponding data sets. In practice, users often apply the “Haldane-Anscombe correction” (Haldane, , 1956; Anscombe, , 1956) to handle sampling zeros in the OR estimation. This correction consists of adding an offset of $+0.5$ to all four cells in the contingency table to avoid issues in computation.

### v2

@@ 3.1.3 Check if data matches scope of study
If method failure persists after ruling out an incorrect use in the execution stage, authors should assess whether real-world users are likely to encounter such data (and, therefore, this method failure). This assessment is shaped by the study’s scope and is naturally not purely objective. Importantly, associated considerations should be reported in the study since they carry valuable information and affect the interpretation of the study’s results.
Consider the case of method failure caused by the method’s inability to make required calculations due to an “extreme” structure of the given simulated data set. For instance, White et al., (2024) conduct an illustrative simulation study that examines methods for handling missing values in confounders in epidemiological data. They observe failure of an imputation method, and investigations of the corresponding data reveal that it occurs when outcome or exposure are highly imbalanced for individuals with an observed confounding variable. From this information, they conclude that “the data-generating mechanism is too extreme and should be changed to generate more outcome events” (White et al., , 2024).
If, given the scope of the study, certain data set characteristics are deemed unrealistic for the targeted readers, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic, in line with White et al., (2024). Similarly, “too extreme” data may also affect real data-based studies. If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely. See Section 4.2 for an illustration. However, if the data sets causing failure are deemed relevant to the (simulation or benchmark) study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation. In particular, they may modify certain parameters of the concerned method(s) (see Section 3.1.4) or resort to a fallback (see Section 3.1.5). To be able to assess whether data structures associated with failure are too extreme, certain considerations should already be made in the planning stage. They should be aware of the scope and goal of their study: is it the goal to examine the general behavior of the method(s) (i.e., earlier stages of research) or to provide concrete recommendations for real-world practitioners (later stages of research)? Is the focus on testing the methods under regular or more extreme circumstances? If aiming to provide recommendations to real-world users, what degree of extremity in the data remains realistic for them?
@@ 3.1.4 Be realistic when modifying method implementations
In earlier stages of research, it may make sense to modify a method’s implementation entirely (i.e., for all simulation repetitions/benchmark data sets) when encountering method failure in the execution stage. For instance, authors may conclude after execution that the default number of maximum likelihood iterations is generally too low, or that the given optimization algorithm may fail frequently, and that an alternative optimizer is more reliable. This way, authors may modify a greater variety of parameters to see when the method is successful and when it fails. Although this early-stage research is not yet directly aimed at real-world users, authors should remember that it will be used by them later. Therefore, certain practical limitations on parameter modifications should apply to ensure that the method remains feasible for future users.
Also, authors should be careful about increasing the runtimes of affected method(s). Especially raising the number of iterations in optimization algorithms can increase runtime or memory consumption, potentially shifting method failure rather than solving it. A special case of modifying a method’s implementation often only arises after the execution stage of the study: investigations following the failure of a method reveal code bugs in the implementation of a method. Especially when their programming experience is high, authors may fix this bug and notify the method developers. Ultimately, the performance assessment of the given method is only meaningful if the authors stress and justify their parameter choices when reporting the study.
@@ 3.1.5 Consider a fallback strategy
When a comparison study aims to provide guidance to real-world users (i.e., in later phases of research), it makes sense to mimic their behavior when encountering method failure in the execution stage. How would they proceed when their method of choice failed on their data set? Morris et al., (2019) argue that they would usually not give up but naturally resort to an alternative method, a fallback, that successfully provides an output. Therefore, it can make sense to transfer this fallback strategy into the context of method comparison studies and evaluate the performance of method pipelines of the type
“use method A if it successfully produces an output, use method B otherwise.”
@@
Incorporating fallbacks indirectly yields automatic handling of method failure in case they affect method A but not method B, allowing for the comparison of (absolute and relative) unconditional performance when chosen carefully and in a way that there is a suitable fallback for each instance of method failure. Note that when using fallback strategies, we no longer evaluate the “pure” performance of method A. However, it is clear from Section 2.2.1 that this is unattainable in the first place.
The choice of a suitable fallback, again, depends on the scope of the study and the targeted users. Should it serve as a guide for experienced or rather inexperienced users? When targeting inexperienced real-world users, authors may choose a completely different method as fallback. For instance, Zapf et al., (2024) suggest from the results of their study that “the approach by Frömke et al (2022) could be used as a fallback strategy in case of non-convergence [of the overall best-performing method] because it always yields results” (Zapf et al., , 2024), where the approach by Frömke et al., (2022) is a non-iterative procedure, therefore never affected by non-convergence by design. When addressing experienced users, an alternative implementation of the given method (in variation of Section 3.1.4) may make a suitable fallback.
Many considerations and decisions regarding fallback strategies should already be made in the planning stage of a study. This includes, first and foremost, if using fallbacks is suitable to the given study, and based on this decision and the scope of the study, whether a completely new method or a modified implementation should be used as fallback. Authors should further be aware that the neutrality principle of late phase comparison studies should always apply, that is, authors should avoid putting considerably more effort into only a few methods when choosing fallbacks for them. This becomes clear, for instance, when modifying the implementation of a method, as, e.g., parameter modifications can strongly influence performance (Weber et al., , 2019). In the same context, it is important to avoid selective reporting as it can lead to over-optimistic results; that is, authors should disclose all tested pipelines, not just the best ones. It is therefore advisable to reduce the data-driven construction and the complexity of pipelines, which can be achieved by the following considerations. First, authors should decide on a maximum number of fallbacks that can be considered to realistically reflect the behavior of real users (as they are unlikely to try out an indefinite number of methods). Suitable combinations of methods and fallbacks may be preliminarily constructed based on the similarity of ease of use (e.g., a user is unlikely to resort to a fallback that is considerably more difficult to apply than the original method). Alternatively, inspired by Zapf et al., (2024), it may make sense to define a default fallback to all instances of method failure if a method has proven to be particularly computationally robust in previous comparison studies.
Finally, authors should be aware that instances of all fallbacks failing on a given data set cannot always be avoided, even when using a computationally robust “default” fallback. Many fallbacks (and generally many methods) failing on the same data set may indicate that it is particularly “challenging.” Authors should then investigate the data set closely to check whether it suits the scope of the study (see Section 3.1.3) and consider removing it if it is deemed unrealistic for real users. Otherwise, reporting the corresponding characteristics of the data set alongside the methods that successfully produce an output may provide valuable insights to real users.
3.1.6 Special case: Runtime issues
In contrast to non-convergence and memory issues, method failure is actively declared by the authors when runtime issues occur in the execution stage. That is, running the methods under investigation on hundreds or thousands of data sets requires time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output. On the other hand, it can sometimes be safely said that real (targeted) users, running the method on a single data set only, are likely to obtain an output. In these cases, when method failure leads to missing rather than undefined performance values, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”). To ensure that this “failure” is independent of data characteristics, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes. Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets in already the planning stage.
@@ 3.1.6 Special case: Runtime issues
3.1.7 Report failure proportions, underlying interplays, applied approaches, and share code
@@
Authors of comparison studies should report any occurrences and proportions of method failure. This can serve as an important criterion for real-world users selecting a suitable method. Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited. Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Zapf et al., (2024) for an example.
@@
Additionally, the interplay of method and data causing failure should be reported. Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” In addition to reporting the corresponding failure proportions, readers can learn in which settings a method is applicable, and no less importantly, in which it is not. Additionally, these insights serve as justification of the chosen handlings to method failure. An example of a corresponding reporting can be found in Hornung et al., (2024), who provide detailed descriptions in the supplement of their paper. Authors of comparison studies should additionally report their applied handlings. Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail. This includes removing any misused methods, removing data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and incorporating fallback strategies. Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
The complexity of the occurrence and handling of method failure, which becomes clear from the previous sections, emphasizes all the more the value of sharing code, data, and intermediate results. Allowing authors of other, “new” comparison studies in-depth investigations to the specific instances of method failure in the given study may help them to a greater understanding of this topic–and enhance fruitful discussions on suitable handlings.
3.1.8 Finally: think ahead and pre-register, but remain flexible
@@
The previous examples and those that will be outlined in Section 4 make clear that method failure and its cause(s), related to complex dynamics between factors such as method and data set, are often difficult to anticipate. While pre-specifying the handling of undefined values has the advantage of reducing analytical uncertainty and its dangers in terms of selective reporting (Siepe et al., , 2024), pre-specifying a standard procedure before conducting the study may sometimes be delicate. Making method failure less unexpected (and pre-specifying a procedure for handling method failure less hazardous) can be achieved through pilot experiments. They test whether an experimental design is appropriate for the specific context in which it is applied, and are recommended in several textbooks on Design and Analysis of Experiments (Cohen, , 1995; Dean et al., , 2017; Montgomery, , 2020). For example, a pilot experiment might use a fraction of the available data sets (e.g., $20\%$) in the comparison study to test whether the specified method implementations produce reasonable results. The remaining $80\%$ is used for the main experiment, which may be refined based on what is learned in the pilot. Pilot studies can also help authors anticipate possible method failures for which they can then pre-specify adequate handlings. Note that it is important to clearly report the results of the pilot study.
Despite pre-registration, researchers should be aware that unforeseen issues may appear and require adaptations of the pre-specified strategy (see Lakens, (2024) for a discussion of the consequences of deviations from pre-registration on the severity and validity of inferences). As a result, error-handling mechanisms should be included in the code right from the beginning, including the storage of error messages (White et al., , 2024).
3.2 Slimmed-down recommendations for the busy researcher (looking to do just enough)
@@
For some authors, there may be time constraints that do not allow for any (in-depth) implementations of our recommendations from the previous Section 3.1, however, authors nevertheless wish to aggregate results over data sets despite the existence of undefined values caused by method failure. While we recommend implementing our in-depth recommendations as comprehensively as possible, we suggest the following slimmed-down analysis and reporting strategy when time and resources are limited. Authors may present performance results after (i) discarding the corresponding data sets for the failing methods, (ii) discarding these data sets for all methods, and (iii) report the failure proportions for all methods, ideally as an additional performance measure. This three-fold analysis and reporting enables more differentiated assessments of the methods’ performances and avoids the preference of methods based solely on how data sets associated with method failure are discarded. It also reveals how successful methods perform on data sets for which others fail (and how this influences the method comparison). When the graphical comparison involves boxplots of performance values, authors could indicate method failure (and the corresponding frequency) as scatters of contrasting color above or below the box. For instance, for bias, they could be indicated above the plot (as a value higher than all biases observed in the study), whereas for accuracy, it could be indicated below the plot.
@@
However, following Section 2.2.1, authors should be aware and clearly state in their manuscript that any form of aggregating performance results despite existing undefined values does not allow an assessment of unconditional performance of all methods under investigation.
@@ 4 Illustrations
In the following, we illustrate the guidelines presented in Section 3 through two fictive comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling.
Note that our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison rather than the method comparison itself. This means that both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study, exclusively focusing on those aspects of the studies that are relevant to the issue of method failure. Importantly, our goal is not to provide insights into the actual performance of the compared methods.
@@ 4.1.1 General
This fictive statistical simulation study deals with the comparison of odds ratio estimation methods. Outcome and exposure are binary ($X,Y\in\left\{0,1\right\}$) and the observed number of subjects with $X=i$ and $Y=j$ is denoted as $n_{ij}$. The scenarios are defined based on different sample sizes $n.obs$, underlying true OR values, and probabilities of exposure $p_{x}$. Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations ($n_{ij}=0$). These “sampling zeros” cause calculation issues (e.g., division by 00) for many OR estimation methods, leading to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$” for the corresponding data sets. In practice, users often apply the “Haldane-Anscombe correction” (Haldane, , 1956; Anscombe, , 1956) to handle sampling zeros in the OR estimation. This correction consists of adding an offset of $+0.5$ to all four cells in the contingency table to avoid issues in computation.